# MaxSup: Overcoming Representation Collapse in Label Smoothing

### v3

@@ 1 Introduction
Multi-class classification [lecun1998mnist, russakovsky2015imagenet] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal. In practice, however, classes often share low-level features [silla2011survey, zeiler2014visualizing] or exhibit high-level semantic similarities [chen2021hsva, novack2023chils, yi2022exploring], rendering the one-hot assumption overly restrictive. Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [guo2020online].
To moderate overconfidence, szegedy2016rethinking introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class. LS has since become prevalent in image recognition [he2016deep, liu2021swin, touvron2021training, zhou2022sp] and neural machine translation [alves2023steering, gao2020towards], often boosting accuracy and calibration [muller2019does]. Yet subsequent work indicates that LS can overly compress features into tight clusters [kornblith2021better, sariyildiz2022no, xu2023quantifying], hindering intra-class variability and transferability [feng2021rethinking]. In parallel, zhu2022rethinking found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
In this paper, we reveal that LS’s training objective inherently contains an error amplification term. This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section˜3.1, Table˜1). Building on zhu2022rethinking, we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics. Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table˜2), a phenomenon corroborated by Grad-CAM visualizations (Figure˜2).
@@ 2.1 Regularization
Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity. Classical methods like $\ell_{2}$ [krogh1991simple] and $\ell_{1}$ [zou2005regularization] impose direct penalties on large or sparse weights, while Dropout [Srivastava2014] randomly deactivates neurons to discourage over-adaptation. In the realm of loss-based strategies, Label Smoothing (LS) [szegedy2016rethinking] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [muller2019does]. Variants such as Online Label Smoothing (OLS) [zhang2021delving] and Zipf Label Smoothing (Zipf-LS) [liang2022efficient] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions. However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section˜3.1, Table 1). Other loss-based regularizers focus on alternative aspects of the predictive distribution. Confidence Penalty [pereyra2017regularizing] penalizes the model’s confidence directly, while Logit Penalty [dauphin2021deconstructing] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [kornblith2021better]. Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section˜4.1). Unlike the aforementioned methods, MaxSup enforces regularization by penalizing only the top-1 logit $z_{\textit{max}}$ rather than the ground-truth logit $z_{gt}$. In LS-based approaches, suppressing $z_{gt}$ for misclassified samples can worsen errors, whereas MaxSup applies a uniform penalty regardless of whether the model’s prediction is correct. Consequently, MaxSup avoids the error amplification effect, retains richer intra-class diversity (see Table 2), and achieves robust transfer performance across diverse datasets and model families (see Table 3).
@@ 2.2 Studies on Label Smoothing
Label Smoothing has also been studied extensively under knowledge distillation. For instance, yuan2020revisiting observed that LS can approximate the effect of a teacher–student framework, while shen2021label investigated its role in such pipelines more systematically. Additionally, chandrasegaran2022revisiting demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes. Concurrently, kornblith2021better showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance. From a Neural Collapse perspective [zhou2022all, guo2024cross], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in xu2023quantifying. Our goal is to overcome LS’s inherent error amplification effect. Rather than adjusting how the smoothed label distribution is constructed (as in OLS or Zipf-LS), MaxSup directly penalizes the highest logit $z_{\textit{max}}$. This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section˜3.2). Moreover, MaxSup integrates seamlessly into standard training pipelines, introducing negligible computational overhead beyond substituting the LS term.
@@ 3.1 Revisiting Label Smoothing
These issues with LS on misclassified samples have also been systematically observed in prior work [xia2024understanding]. By precisely disentangling these two components (regularization vs. error amplification), we can design a more targeted and effective solution.
@@
Ablation Study on LS Components. To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [touvron2021training] without Mixup or CutMix. As indicated in Table˜1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy. In contrast, our MaxSup omits the error amplification component and leverages only the beneficial regularization, thereby boosting accuracy beyond that of standard LS. Specifically, Table˜1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%). Disabling only the error amplification while retaining the regularization yields a slight but measurable improvement (75.98% vs. 75.91%). By fully removing error amplification and faithfully preserving the helpful aspects of LS, our MaxSup achieves 76.12% accuracy, clearly and consistently outperforming LS. This result underscores that MaxSup directly tackles LS’s fundamental shortcoming by maintaining a consistent and meaningful regularization signal—even when the top-1 prediction is incorrect.
@@ 3.2 Max Suppression Regularization
Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl(1-\tfrac{1}{K}\bigr)$, while all other logits slightly increase by $\tfrac{\alpha}{K}$. In misclassified cases, the ground-truth logit $z_{gt}$ is spared from penalization, avoiding the error-amplification issue seen in LS. For completeness, Appendix B provides the full gradient derivation. While [xia2024understanding] conducted a related gradient analysis of the training loss, it focuses specifically on the setting of selective classification, and examines a posthoc logit normalization technique to mitigate confidence calibration issues. However, this approach addresses only the overconfidence problem of label smoothing (LS), without tackling representation collapse. Moreover, our work presents a logit-level reformulation of LS that provides a deeper theoretical understanding of why LS amplifies errors.
@@ 4.1.1 Intra-Class Variation and Transferability
Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[kornblith2021better] for ResNet-50 trained on ImageNet-1K. Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
@@ 4.1.2 Connection to Logit Penalty
As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [kornblith2021better]. Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
@@ 4.2.1 Experiment Setup
Model Training Configurations.We evaluate both convolutional (ResNet[he2016deep], MobileNetV2 [sandler2018mobilenetv2]) and transformer (DeiT-Small [touvron2021training]) architectures on ImageNet [krizhevsky2012imagenet]. For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048. The initial learning rate is 0.85 and is annealed via a cosine schedule. [^2] Additional training hyperparameters follow the FFCV scripts at https://github.com/libffcv/ffcv. See Appendix F for further details. We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$. For DeiT-Small, we use the official codebase [touvron2021training], training from scratch without knowledge distillation to isolate MaxSup’s contribution. CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[liang2022efficient] and Online Label Smoothing [zhang2021delving]. Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper. Except for any method-specific hyperparameters, all other core training settings remain identical to the baselines. Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix˜G). This ensures a fair comparison under consistent and reproducible training protocols.
@@ 4.2.2 Experiment Results
ConvNet Comparison. Table4 shows results for MaxSup alongside various label-smoothing and self-distillation methods on both ImageNet and CIFAR-100 benchmarks. Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches. By contrast, OLS [zhang2021delving] and Zipf-LS [liang2022efficient] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
@@
Fine-Grained Classification. Beyond large-scale benchmarks like ImageNet, we further evaluate MaxSup on two fine-grained visual recognition tasks: CUB-200-2011 [cub] and Stanford Cars [car]. These datasets pose unique challenges due to subtle inter-class differences, which often expose the limitations of standard regularization approaches. As shown in Table˜6, MaxSup achieves the best performance across both datasets, surpassing LS and its recent variants. This demonstrates that MaxSup encourages the model to learn more discriminative and semantically rich representations that better capture fine-grained attributes, such as textures and part-level details. The consistent improvements on these benchmarks further validate MaxSup’s capacity to generalize across different visual domains and its potential to enhance robustness in recognition scenarios where nuanced feature understanding is critical.
@@
Long-Tailed Classification. To assess the effectiveness of MaxSup under data imbalance, we performed experiments on the CIFAR-10-LT dataset with imbalance ratios of 50 and 100, following the experimental settings described in [tang2020long]. The corresponding results are summarized in Table˜7. The evaluation compares three setups: Focal Loss, Focal Loss + LS, and Focal Loss + MaxSup. Across all imbalance ratios and splits (val/test), MaxSup consistently outperforms both the baseline and LS in overall accuracy, which jointly reflects the many-shot, medium-shot, and low-shot (minor class) performance. For example, at an imbalance ratio of 50 on the test split, MaxSup achieves $81.4\%$ accuracy, outperforming Focal Loss ($76.8\%$) by 4.6 percentage points, and LS ($80.5\%$) by 0.9 percentage points. These results indicate that MaxSup achieves a better trade-off between many- and medium-shot accuracy. While it does not fully resolve the challenge of imbalanced classification—especially for minority classes—it shows positive effects and offers a promising direction for further extension.
@@
Corrupted Image Classification To evaluate the effectiveness of MaxSup on out-of-distribution (OOD) settings, we also conducted experiments on CIFAR10-C benchmark [hendrycksbenchmarking] shown in Table˜8 following settings in [heinonen2025robust]. Table˜8 reports the performance of MaxSup and Label Smoothing (LS) on this benchmark using ResNet-50 as the backbone. Specifically, LS yields a better NLL (1.5730 vs. 1.8431), implying more confident probabilistic predictions. However, MaxSup achieves a better ECE (0.1479 vs. 0.1741), indicating better calibration of the predicted confidence scores. These results validate that MaxSup remains effective on OOD datasets, achieving performance comparable to LS across all three metrics.
@@ 4.3 Evaluation on Semantic Segmentation
We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework. [^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [xiao2018unified] architecture with a DeiT-Small backbone, trained on ADE20K. Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section˜4.2.2).
@@
Figure 2: Grad-CAM [Selvaraju_2019] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row). The first row shows the original images. Compared to Label Smoothing, MaxSup more effectively filters out non-target regions and highlights essential features of the target class, reducing instances where the model partially or entirely focuses on irrelevant areas.

### v4

@@ 1 Introduction
Multi-class classification [19, 26] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal. In practice, however, classes often share low-level features [31, 44] or exhibit high-level semantic similarities [3, 24, 42], rendering the one-hot assumption overly restrictive. Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [9].
To moderate overconfidence, Szegedy et al. [34] introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class. LS has since become prevalent in image recognition [10, 22, 36, 47] and neural machine translation [1, 6], often boosting accuracy and calibration [23]. Yet subsequent work indicates that LS can overly compress features into tight clusters [15, 28, 41], hindering intra-class variability and transferability [5]. In parallel, Zhu et al. [48] found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
In this paper, we reveal that LS’s training objective inherently contains an error amplification term. This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section˜3.1, Table˜1). Building on Zhu et al. [48], we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics. Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table˜2), a phenomenon corroborated by Grad-CAM visualizations (Figure˜2).
@@ 2.1 Regularization
Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity. Classical methods like $\ell_{2}$ [18] and $\ell_{1}$ [49] impose direct penalties on large or sparse weights, while Dropout [32] randomly deactivates neurons to discourage over-adaptation. In the realm of loss-based strategies, Label Smoothing (LS) [34] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [23]. Variants such as Online Label Smoothing (OLS) [45] and Zipf Label Smoothing (Zipf-LS) [21] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions. However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section˜3.1, Table 1). Other loss-based regularizers focus on alternative aspects of the predictive distribution. Confidence Penalty [25] penalizes the model’s confidence directly, while Logit Penalty [4] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [15]. Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section˜4.1). Unlike the aforementioned methods, MaxSup enforces regularization by penalizing only the top-1 logit $z_{\textit{max}}$ rather than the ground-truth logit $z_{gt}$. In LS-based approaches, suppressing $z_{gt}$ for misclassified samples can worsen errors, whereas MaxSup applies a uniform penalty regardless of whether the model’s prediction is correct. Consequently, MaxSup avoids the error amplification effect, retains richer intra-class diversity (see Table 2), and achieves robust transfer performance across diverse datasets and model families (see Table 3).
@@ 2.2 Studies on Label Smoothing
Label Smoothing has also been studied extensively under knowledge distillation. For instance, Yuan et al. [43] observed that LS can approximate the effect of a teacher–student framework, while Shen et al. [30] investigated its role in such pipelines more systematically. Additionally, Chandrasegaran et al. [2] demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes. Concurrently, Kornblith et al. [15] showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance. From a Neural Collapse perspective [46, 8], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in Xu and Liu [41]. Our goal is to overcome LS’s inherent error amplification effect. Rather than adjusting how the smoothed label distribution is constructed (as in OLS or Zipf-LS), MaxSup directly penalizes the highest logit $z_{\textit{max}}$. This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section˜3.2). Moreover, MaxSup integrates seamlessly into standard training pipelines, introducing negligible computational overhead beyond substituting the LS term.
@@ 3.1 Revisiting Label Smoothing
These issues with LS on misclassified samples have also been systematically observed in prior work [39]. By precisely disentangling these two components (regularization vs. error amplification), we can design a more targeted and effective solution.
@@
Ablation Study on LS Components. To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [36] without Mixup or CutMix. As indicated in Table˜1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy. In contrast, our MaxSup omits the error amplification component and leverages only the beneficial regularization, thereby boosting accuracy beyond that of standard LS. Specifically, Table˜1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%). Disabling only the error amplification while retaining the regularization yields a slight but measurable improvement (75.98% vs. 75.91%). By fully removing error amplification and faithfully preserving the helpful aspects of LS, our MaxSup achieves 76.12% accuracy, clearly and consistently outperforming LS. This result underscores that MaxSup directly tackles LS’s fundamental shortcoming by maintaining a consistent and meaningful regularization signal—even when the top-1 prediction is incorrect.
@@ 3.2 Max Suppression Regularization
Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl(1-\tfrac{1}{K}\bigr)$, while all other logits slightly increase by $\tfrac{\alpha}{K}$. In misclassified cases, the ground-truth logit $z_{gt}$ is spared from penalization, avoiding the error-amplification issue seen in LS. For completeness, Appendix B provides the full gradient derivation. While [39] conducted a related gradient analysis of the training loss, it focuses specifically on the setting of selective classification, and examines a posthoc logit normalization technique to mitigate confidence calibration issues. However, this approach addresses only the overconfidence problem of label smoothing (LS), without tackling representation collapse. Moreover, our work presents a logit-level reformulation of LS that provides a deeper theoretical understanding of why LS amplifies errors.
@@ 4.1.1 Intra-Class Variation and Transferability
Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[15] for ResNet-50 trained on ImageNet-1K. Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
@@ 4.1.2 Connection to Logit Penalty
As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [15]. Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
@@ 4.2.1 Experiment Setup
Model Training Configurations.We evaluate both convolutional (ResNet[10], MobileNetV2 [27]) and transformer (DeiT-Small [36]) architectures on ImageNet [17]. For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048. The initial learning rate is 0.85 and is annealed via a cosine schedule. [^2] Additional training hyperparameters follow the FFCV scripts at https://github.com/libffcv/ffcv. See Appendix F for further details. We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$. For DeiT-Small, we use the official codebase [36], training from scratch without knowledge distillation to isolate MaxSup’s contribution. CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[21] and Online Label Smoothing [45]. Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper. Except for any method-specific hyperparameters, all other core training settings remain identical to the baselines. Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix˜G). This ensures a fair comparison under consistent and reproducible training protocols.
@@ 4.2.2 Experiment Results
ConvNet Comparison. Table4 shows results for MaxSup alongside various label-smoothing and self-distillation methods on both ImageNet and CIFAR-100 benchmarks. Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches. By contrast, OLS [45] and Zipf-LS [21] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
@@
Fine-Grained Classification. Beyond large-scale benchmarks like ImageNet, we further evaluate MaxSup on two fine-grained visual recognition tasks: CUB-200-2011 [37] and Stanford Cars [16]. These datasets pose unique challenges due to subtle inter-class differences, which often expose the limitations of standard regularization approaches. As shown in Table˜6, MaxSup achieves the best performance across both datasets, surpassing LS and its recent variants. This demonstrates that MaxSup encourages the model to learn more discriminative and semantically rich representations that better capture fine-grained attributes, such as textures and part-level details. The consistent improvements on these benchmarks further validate MaxSup’s capacity to generalize across different visual domains and its potential to enhance robustness in recognition scenarios where nuanced feature understanding is critical.
@@
Long-Tailed Classification. To assess the effectiveness of MaxSup under data imbalance, we performed experiments on the CIFAR-10-LT dataset with imbalance ratios of 50 and 100, following the experimental settings described in [35]. The corresponding results are summarized in Table˜7. The evaluation compares three setups: Focal Loss, Focal Loss + LS, and Focal Loss + MaxSup. Across all imbalance ratios and splits (val/test), MaxSup consistently outperforms both the baseline and LS in overall accuracy, which jointly reflects the many-shot, medium-shot, and low-shot (minor class) performance. For example, at an imbalance ratio of 50 on the test split, MaxSup achieves $81.4\%$ accuracy, outperforming Focal Loss ($76.8\%$) by 4.6 percentage points, and LS ($80.5\%$) by 0.9 percentage points. These results indicate that MaxSup achieves a better trade-off between many- and medium-shot accuracy. While it does not fully resolve the challenge of imbalanced classification—especially for minority classes—it shows positive effects and offers a promising direction for further extension.
@@
Corrupted Image Classification To evaluate the effectiveness of MaxSup on out-of-distribution (OOD) settings, we also conducted experiments on CIFAR10-C benchmark [12] shown in Table˜8 following settings in [11]. Table˜8 reports the performance of MaxSup and Label Smoothing (LS) on this benchmark using ResNet-50 as the backbone. Specifically, LS yields a better NLL (1.5730 vs. 1.8431), implying more confident probabilistic predictions. However, MaxSup achieves a better ECE (0.1479 vs. 0.1741), indicating better calibration of the predicted confidence scores. These results validate that MaxSup remains effective on OOD datasets, achieving performance comparable to LS across all three metrics.
@@ 4.3 Evaluation on Semantic Segmentation
We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework. [^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [40] architecture with a DeiT-Small backbone, trained on ADE20K. Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section˜4.2.2).
@@
Figure 2: Grad-CAM [29] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row). The first row shows the original images. Compared to Label Smoothing, MaxSup more effectively filters out non-target regions and highlights essential features of the target class, reducing instances where the model partially or entirely focuses on irrelevant areas.